OpenStudy (anonymous):

write 68 billion in scientific notation

jimthompson5910 (jim_thompson5910):

"billion" means "10^9" saying "68 billion" means \(\Large 68 \times 10^9\) now convert to scientific notation (fill in the blank and the empty box below) \[\Large 68 \times 10^{9} = \underline{ \ \ \ \ \ \ } \times 10^{{}^{\Huge \boxed{}}}\]

OpenStudy (anonymous):

6.8 x 10 to the 10th

jimthompson5910 (jim_thompson5910):

very good

OpenStudy (anonymous):

\[6.8 x 10^{10}\]